Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, December 04, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 04, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/04/98 (Friday)

Segment 1 report for 12/4/98: Backfill began at the segment that was tied-in on the north end of the Cuddy Creek directional drill (EMP 25.12) . On Lebec Rd. excavation through the road continued to 21.54 and a segment of pipe was tied-in at 21.50. On Tejon Bench the backfill crew was unable to work due to the muddy conditions from rain on the previous days. Welding, x-ray and coating occurred 16.59 to 16.55. On Grapevine road, bore equipment was being delivered to the location where the I-5 northbound lane bore is at EMP 16.49. Welding, x-ray and coating continued from EMP's 15.76 to 15.71. Electrical work continued at Grapevine Pump Station. Boring at the Stream Crossing #9 bore continued (EMP 15.35). At the Hwy. 166 bore (EMP 1.18) backfill began. Work on the office interior continued at Emidio Pump Station.

Segment 2B (Castaic to Sylmar): Report for 12-4-98. Clean up of the directional drill site continued at the Santa Clara river and Feedmill Rd. Fiber optic crews pulled 100 strand cable between Pico Cyn Rd and the I-5 overcrossing at EMP 69.99. Meanwhile, trenching crews have completed ditch on segment 2B completing work at EMP 71.4. In addition, pipe and welding crews have one tie-in to complete at 71.4, near Weldon Cyn. X-ray and coating crews are working near EMP 72.15 while 600' of slurry backfill and fiber optic duct were completed to EMP 72.23. Asphalt was completed to EMP 72.36 and crews are removing spoils from Weldon Cyn Rd. Near Sagecrest Circ, crews are excavating the roadway to repair a section of damaged 20" pipe.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 4 December 1998: 3A: The bore crew finally freed their augers from under Tuxford Street. The crew began tearing down and loading up their equipment, this being the last bore on Segment 3. Backfill spent a long day continuing their work at Tujunga Wash. Welding, x-ray, and coating continued their work at Pacoima Wash Bridge crossing, focusing on the south side tie in. Excavation crews dug ditch on the north side of the bridge. GPS survey and soil compaction were done today. 3B: NIGHT work consisted of electrical conduit work at the end of Cornfield Yard, across North Spring Street. 3C: Backfill crews were the sole crew on Segment 3C today. Their work focused on the area directly adjacent to the aqueduct crossing.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page